OpenStudy (anonymous):

I am doing a test and want to make sure that I have this answer correct so can someone please check it for me.

OpenStudy (anonymous):

\[\sqrt[7]{128(x-1)^7} = 6(x-1)\]

OpenStudy (anonymous):

Is that correct

OpenStudy (anonymous):

no its not why 6?

OpenStudy (anonymous):

\[6^{7} = 128\]

OpenStudy (anonymous):

no \[2^{7}=128\]

OpenStudy (anonymous):

\[6^{7}=6*6*6*6*6*6*6\] which is LOT more than 128

OpenStudy (anonymous):

sorry 2 not 6

OpenStudy (anonymous):

sorry I had it written down right just typed it wrong

OpenStudy (anonymous):

ok so if you have 2 its correct

OpenStudy (anonymous):

but would it be right as 2(x-1)

OpenStudy (anonymous):

yes 2(x-1) is right

OpenStudy (anonymous):

ok thanks and sorry about the confusion

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

my stupid but still put 6 on my freaking test